# A Generative Approach to Credit Prediction with Learnable Prompts for Multi-scale Temporal Representation Learning

### v3

@@ Abstract
Recent industrial credit scoring models remain heavily reliant on manually tuned statistical learning methods. While deep learning offers promising solutions, its effectiveness is often limited by the complexity of financial data, particularly in long-horizon scenarios. In this work, we propose FinLangNet [^1] https://github.com/leiyu0210/FinLangNet, which addresses credit scoring by reframing it as the task of generating multi-scale distributions of a user’s future behavior. Within this framework, tabular data is transformed into sequential representations, enabling the generation of user embeddings across multiple temporal scales. Inspired by the recent success of prompt-based training in Large Language Models (LLMs), FinLangNet also introduces two types of prompts to model and capture user behavior at both the feature-granularity and user-granularity levels. Experimental results demonstrate that FinLangNet outperforms the online XGBoost benchmark, achieving a 7.2% improvement in KS metric performance and a 9.9% reduction in the relative bad debt rate. Furthermore, FinLangNet exhibits superior performance on public UEA archives, underscoring its scalability and adaptability in time series classification tasks.
@@ 4.2 Non-Sequential Module
Table 1: Performance Metrics for Multiple Models across Multiple Labels. The models in the upper part of the table use tabular data for modeling (internal features are aligned), and the models in the lower part use time series and non-time series features (internal features are also aligned). Other training and evaluation parameters remain consistent.
@@ 4.3.2 Loss Function
5 Experiment
5.1 Experiment Setup
Initially, our experiments are carried out on real business data. We sampled data from over 7 million users, resulting in a dataset of approximately 2 million samples. The dataset contains 6 evaluation tasks, representing overdue status at different time scales. The evaluation indicators of each task are KS (Massey Jr, 1951) and AUC. We compared the online benchmark model and some deep model methods under the same training parameter conditions. Additional details of our dataset are provided in Appendix Table 5. Detailed training parameters can be found in Appendix Table 7.
5.2 Performance on Credit Prediction
As demonstrated in Table 1, we evaluated the performance of our proposed model across six different time scales, denoted as $\tau$ ranging from 1 to 6. For each corresponding time scale, we labeled the results as $y_{1}$ through $y_{6}$. To benchmark our approach, we compared it against several models, including XGBoost (online baseline models), MLP, and graph-based models such as GraphSAGE and ChebConv. Our model was trained just once and was able to generalize across all six time scales without the need for retraining. Our evaluation was based on key metrics such as AUC and KS. As shown in Table 1, when dealing with tabular data in the risk control domain, XGBoost remains one of the most competitive methods compared to other deep learning approaches. The results clearly demonstrate that our model outperformed all baseline approaches across every individual time scale. On average, compared with online baseline models, our model achieved a 7.2% improvement. This significant improvement highlights the robustness and scalability of our approach.
Moreover, given that our experiment dataset is composed of both sequential and non-sequential data, with the majority being sequential, we conducted additional ablation studies on the sequence modeling component. Specifically, we replaced our sequence module with several other models, including LSTM, GRU, Mamba, Transformer, and TimesNet. The results in Table 1 indicate that our model consistently outperformed all of these alternatives, demonstrating its effectiveness in handling sequential data.
In summary, the experimental results validate the superiority of our model in both average performance and adaptability across multiple time scales. Furthermore, its modularity and efficiency make it highly scalable for various downstream tasks, especially in scenarios with heavy sequential data dependencies.
@@ 5.3.1 Impact of Sequential Data Length
Considering that different users have varying sequence lengths ($T$), it is sometimes necessary to perform sequence truncation or segmentation to handle these variations. To investigate the impact of sequence length on model performance, we conducted ablation experiments where only the length of the sequential data was varied, while keeping non-sequential data constant. As shown in Table 2, the full-length data yielded an average KS value of 33.71%. This decreased marginally to 33.62% at half-length and further to 33.49% at a quarter-length.
Table 3: Influence of Different Modules in Ablation Studies on the $y_{1}(\tau=1)$
Configuration KS AUC
Baseline (No Prompts) 0.3262 0.7254
+ Prompt ($\widetilde{\phi}_{c}$) 0.3279 0.7265
+ Prompt ($P_{s}$) 0.3299 0.7278
+ Both Prompts ($\widetilde{\phi}_{c}$ + $P_{s}$) 0.3408 0.7355
@@ 5.3.2 Role of Different Module
The second area of our ablation study, as shown in Table 3, aims to investigate the contributions of the two prompts introduced in SRG (our sequence module). Specifically, the first prompt is the feature granularity prompt $\widetilde{\phi}_{c}$, which captures domain-specific feature-level contextual information, while the second prompt is the data-source granularity prompt $P_{s}$, designed to encode interactions and patterns at the data-source level.
Our results show that both prompts contribute positively to model performance, as evidenced by consistent improvements across key evaluation metrics when either prompt is introduced. However, a closer examination reveals that the data-source granularity prompt $P_{s}$ provides a more significant performance gain compared to the feature granularity prompt $\widetilde{\phi}_{c}$. This result suggests that capturing relationships and dependencies at the data-source level is particularly crucial for improving the model’s understanding and utilization of sequential data.
@@
In summary, these findings highlight the importance of incorporating both feature-level and data-source-level contextual information into the model. While each type of prompt contributes independently to the performance enhancement, the greater impact of $P_{s}$ underscores the value of effectively leveraging data-source granularity in sequential modeling. These insights further solidify the advantages of SRG’s design for robust and interpretable sequence prediction tasks.

### v4

@@ Abstract
Recent industrial credit scoring models remain heavily reliant on manually tuned statistical learning methods. While deep learning offers promising solutions, its effectiveness is often limited by the complexity of financial data, particularly in long-horizon scenarios. In this work, we propose FinLangNet [^1] https://github.com/didiglobal-fintech-credit-risk/FinLangNet, which addresses credit scoring by reframing it as the task of generating multi-scale distributions of a user’s future behavior. Within this framework, tabular data is transformed into sequential representations, enabling the generation of user embeddings across multiple temporal scales. Inspired by the recent success of prompt-based training in Large Language Models (LLMs), FinLangNet also introduces two types of prompts to model and capture user behavior at both the feature-granularity and user-granularity levels. Experimental results demonstrate that FinLangNet outperforms the online XGBoost benchmark, achieving a 7.2% improvement in KS metric performance and a 9.9% reduction in the relative bad debt rate. Furthermore, FinLangNet exhibits superior performance on public UEA archives, underscoring its scalability and adaptability in time series classification tasks.
@@ 2 Related Work
Meanwhile, we also investigated several LLM-based approaches. In real-world applications, general-purpose large language models often fall short in meeting specific business needs. For example, although CALM(Feng et al., 2023) is a financial-domain LLM, it is not tailored for credit scoring tasks. CALM outputs binary classifications (e.g., Good/Bad), while our model, FinLangNet, produces continuous delinquency probabilities, which can be directly mapped to credit scores. These scores are crucial for making fine-grained decisions such as customer acceptance or rejection based on cutoff thresholds. Moreover, model size and output stability are also critical considerations in production environments, where large models like CALM may not perform reliably or efficiently.
@@ 4.2 Non-Sequential Module
Table 1: Performance comparison across multiple models and labels. The upper part lists models using tabular features, while the lower part includes models combining sequential and non-sequential features. All models share the same training and evaluation settings.
@@ 4.3.2 Loss Function
5 Experiment
5.1 Experiment Setup
Initially, our experiments are carried out on real business data. We sampled data from over 7 million users. The dataset contains 6 evaluation tasks, representing overdue status at different time scales. The evaluation indicators of each task are KS (Massey Jr, 1951) and AUC. We compared the online benchmark model and some deep model methods under the same training parameter conditions. Additional details of our dataset are provided in Appendix Table 5. Detailed training parameters can be found in Appendix Table 7.
5.2 Performance on Credit Prediction
As illustrated in Table 1, we evaluate model performance across six temporal scales, denoted as $\tau$ ranging from 1 to 6, with prediction results labeled as $y_{1}$ through $y_{6}$. Each task is assessed using AUC and KS as evaluation metrics.
In the upper part of the table, we compare traditional tabular-based models. XGBoost, the current online baseline, exhibits strong performance across most tasks and remains highly competitive in risk control scenarios. However, our proposed FinLangNet consistently outperforms XGBoost and other deep learning baselines on all six tasks, achieving an average improvement of 7.2% in key metrics. This demonstrates both the robustness and scalability of our framework in handling complex business data.
The lower section of Table 1 focuses on models tailored for sequential inputs, including GRU (Cho et al., 2014), Transformer (Vaswani et al., 2017), Mamba (Gu and Dao, 2023), TimesNet (Wu et al., 2023), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019) and BST (Chen et al., 2019). All models are trained under identical hyperparameter settings. Even in this setting, FinLangNet achieves superior performance with negligible additional computational cost. We attribute the superior performance of our model to its enhanced ability to generalize across heterogeneous channels and data sources, a limitation commonly observed in standard sequential models that often struggle in real-world scenarios despite performing well on academic datasets. To address this, we introduce two tailored prompts that facilitate cross-source information integration. Furthermore, we reformulate credit risk modeling as a multi-scale representation learning task, enabling comprehensive user risk profiling across the full credit lifecycle. This multi-scale design promotes mutual learning across time horizons, contributing to consistently optimal performance across all tasks.
@@ 5.3.1 Impact of Sequential Data Length
To assess the effect of sequence length on model performance, we conducted a set of ablation experiments in which only the length of the sequential inputs was varied, while non-sequential features were held fixed. As reported in Table 2, shortening the sequence led to only marginal performance degradation, suggesting that the model exhibits robustness to sequence length variation. This also underscores the presence of critical information in early parts of the sequence and highlights the model’s ability to generalize under conditions of incomplete or truncated data.
Table 3: Influence of Different Modules on task $y_{1}(\tau=1)$
Configuration KS(%) AUC(%)
Baseline (No Prompts) 32.62 72.54
+ Prompt ($\widetilde{\phi}_{c}$) 32.79 72.65
+ Prompt ($P_{s}$) 32.99 72.78
+ Both Prompts ($\widetilde{\phi}_{c}$ + $P_{s}$) 34.08 73.55
w/o Sequential model ($O_{\text{ns}}$) 12.53 58.62
w/o Non-sequential model ($O_{m}$) 32.75 72.66
@@ 5.3.2 Role of Different Module
We conducted ablation studies to evaluate the contribution of different components in our model, as summarized in Table 3. First, we assess the impact of the two prompts in the SRG module: the feature-granularity prompt $\widetilde{\phi}_{c}$ and the data-source prompt $P_{s}$. Both prompts individually improve performance over the no-prompt baseline, with $P_{s}$ yielding a larger gain, indicating that modeling cross-source interactions is more critical than capturing fine-grained feature context.
@@
Furthermore, we examine the role of the sequential and non-sequential modules. Removing the sequential module results in a substantial performance drop, confirming its critical role in capturing temporal dynamics. In contrast, removing the non-sequential module leads to only a slight decline, implying that sequential information carries the dominant signal for our task.
These results validate the effectiveness of our dual-prompt strategy and modular design, highlighting the importance of incorporating both hierarchical prompts and multi-modal input streams for robust credit risk modeling.
@@ 6 Online A/B Test
For deployment, the model is converted to ONNX format and served on an L20 GPU cluster. Given the low-frequency nature of credit applications, the system operates under moderate QPS ($\sim 100$) and achieves inference latency within 100 ms, meeting real-time business requirements.
To ensure transparency and trustworthiness, we incorporate a lightweight interpretability framework based on the LIME (Local Interpretable Model-agnostic Explanations) (Ribeiro et al., 2016) method. LIME perturbs the input—by masking or shuffling elements in the sequence—and learns a local surrogate model to identify the most influential features driving predictions. This allows us to interpret the behavior of our sequential model in a locally faithful manner.
We apply this mechanism across different data sources to rank feature importance, which serves multiple practical purposes: (i) guiding feature selection during model development, (ii) supporting issue diagnosis after deployment, and (iii) monitoring feature drift in production.